# A hydrodynamic approach to reproduce multiple spinning vortices in horizontally rotating three-dimensional liquid helium-4

### v1

@@ 2.1 The two-fluid model with spin angular momentum conservation
The mathematical expressions for the three assumptions shown in Fig. 1 can be described in a form similar to that of the conventional two-fluid model derived from the $\rm Schr\ddot{o}dinger$ equation. For clarity, we make the following assumptions. (i) It is a mixture of a viscous and an inviscid fluid. (ii) The ratio of helium atoms in the ground state to the total number of atoms follows quantum statistical mechanics for bosons and thus fluctuates around its statistical average, enabling us to assume that the volume of the fluid fluctuates; in other words, weak compressibility can be assumed. (iii) Local interaction forces are exerted between two different types of fluid particles. Moreover, in another recent two-fluid model, the vortex filament model was solved instead of the inviscid equation, in which the rotation of the vortices can be considered. In contrast, we assume spin angular momentum conservation of fluid particles around their axes; thus, the viscosity term is decomposed into three terms in Eq.  (2).
@@
Let us explain the physical significance of each parameter in Eqs.  (3) and (4) with reference to Refs.  [6, 7]. In Eq. (3), $M$ represents the mass of a helium-4 atom, $N$ indicates the total number of helium-4 atoms in the system, $\pi$ represents the circumference of the circle, and $c$ is the speed of sound. $\hbar$ indicates the reduced Planck’s constant and $k_{B}$ represents the Boltzmann constant. The constant values $\mu$, $p_{0}$, and $\Delta$ are given by ${\rm 1.72\times 10^{-24}~{}g}$, ${\rm 2.1\times 10^{-19}~{}gcms^{-1}}$, and ${\rm 8.9~{}K}$, respectively  [25, 26]. Eq. (3) is always true within the scope of this study because it is valid if $T\ll{\rm 93~{}K}$ [25]. In the elementary excitation model, the first term in the parentheses on the right side of Eq. (3) represents the “photon” contribution, and the second term in the same parentheses represents the “roton” contribution; after finding the total pressure from the summation of the photon and roton contributions, the entropy density in Eq. (3) is obtained by the thermodynamic relation, such that the partial derivative of pressure with respect to temperature [6]. Equation (3) represents the relationship between entropy density $\sigma$ and temperature $T$ in a quantum “ideal gas.” To obtain the value of liquid helium-4, we decompose $\sigma$ as $\sigma=C_{e}\sigma_{0}$, where $C_{e}$ indicates the volume ratio of liquid to gaseous helium-4 under an atmospheric pressure of $\rm 1.01325~{}bar$, which is $1.428\times 10^{-3}$ [27]. Under constant temperature, we calculated $\sigma_{0}$ using Eq.  (3) and maintained a constant value for entropy conservation. Meanwhile, in Eq. (4), $\rho_{0}$ and $p_{0}$ represent the fluid density and pressure in the initial state, respectively. $\alpha_{P}$ is the specific heat ratio that affects system incompressibility. $\beta_{P}$ denotes the reference pressure, which typically corresponds to $p_{0}$. Equations (1)–(4) represent the governing equations of the system.
@@ 2.2 Multi-GPU computing for large-scale particle simulations
Figure 2: Snapshots of particle distribution (left part) and the domain decomposition in the case where approximately 19.6 million fluid particles were divided into 32 subdomains, each of which was assigned to a single GPU.
It is imperative to develop dynamic load balancing among the GPUs to exploit the high computational efficiency of multiple GPUs. Various load balancing techniques have been proposed in related fields, including the slice-grid method, which moves the domain boundary to a spatially partitioned region; graph theory-based partitioning (such as k-means), as represented by ParMETIS [50, 51] and Zoltan [52, 53]; and the hierarchically structured tree method, which uses a recursive structure of cells to partition a domain, as exemplified by kd-tree decomposition [54], orthogonal recursive bisection (ORB) [55, 56, 57], and domain decomposition using a hierarchically structured grid with space-filling curves [58]. Recently, we conducted a study on optimal dynamic domain decomposition for particle simulations of short-range interactions compatible with low byte/flops supercomputers, such as GPU supercomputers. Subsequently, we developed a multi-GPU computing framework that enables large-scale particle simulations using SPH. The SPH computational framework [46] developed by the author was used in this study. Based on the nature of the target problem, the 2D slice grid method was used as the domain decomposition method. Figure 2 illustrates a snapshot of the domain decomposition in the case where the 19.5 million fluid particles were divided into 32 subdomains, each of which was assigned to a single GPU to speed up the simulations.
@@ 3 Large-scale simulations of horizontally rotating 3D liquid helium-4
Let us set the outer diameter of a cylindrical container in the horizontal direction to 0.2 cm with a thickness of 0.04 in the vertical direction (z-direction). The resolutions of $(N_{x},N_{y},N_{z})$ are given by (500, 500, 100), where $N_{a}$ $(a=x,y,z)$ indicates the number of particles in the direction in a rectangular domain that covers the cylindrical container. We set the rotational angular velocity around the cylinder axis (z axis) to 5 rad $\cdot s^{-1}$ counterclockwise. Similar to the two-dimensional cases in our previous work, we set the temperature T to 1.6 K and determined the values of $(\rho_{s}/\rho,\rho_{n}/\rho,\kappa)$ with those at T = 1.6 K listed in Ref. [63]. In addition, $\eta_{n}$ was obtained from the relationship of 0.566 $\kappa\rho_{n}$ with reference to [63]. The shear viscosity $\bar{\eta}$ and rotational viscosity $\bar{\eta_{r}}$ have the value of $\eta_{n}$. We set the parameter $C_{\omega}$, which determines the magnitude of the angular velocity $\omega$ and thereby determines the resulting number of vortices, to be 0.016 to maintain a similar degree of magnitude of the rotational forces as those in 2D cases reported in our previous studies. Normal fluid particles were placed along the inner circumference of the container as walls, and no-slip conditions were imposed on the wall particles as boundary conditions. Conversely, we set slip conditions at the top and bottom of the vessel. In the initial state, superfluid or normal fluid particles were randomly generated in proportion to the density ratio and distributed in the vessel. We set dt to $5.0\times 10^{-6}$ s and computed 180000 time steps, which is 0.9 s in physical time. In the simulation, under the aforementioned conditions, the total number of normal and inviscid fluid particles, including the wall particles, was 19636400. We boosted our simulation using 32 GPUs installed on the Wisteria/BDEC-01 supercomputer at The University of Tokyo, JAPAN. We applied a two-dimensional slice-grid technique to the dynamic domain decomposition in the horizontal direction and decomposed the computational domain into 32 subdomains (four in the horizontal direction and eight in the depth direction), and a single GPU was assigned to each subdomain. Consequently, approximately 5.5 days were required to compute all 180000 time steps of the simulation. For comparison, we performed simulations with a lower resolution using $(N_{x},N_{y},N_{z})$ = (250, 250, 50) to investigate the dependence of the observed phenomena on the computational resolution.

### v2

@@ 2.1 The two-fluid model with spin angular momentum conservation
In this manner, we adopted the following assumptions. (i) It is a mixture of a viscous and an inviscid fluid at finite temperature. (ii) The ratio of helium atoms in the ground state to the total number of atoms follows quantum statistical mechanics for bosons and thus fluctuates around its statistical average, enabling us to assume that the volume of the fluid fluctuates; in other words, weak compressibility can be assumed. (iii) Local interaction forces are exerted between two different types of fluid particles. Moreover, in another recent two-fluid model, the vortex filament model [24, 25, 26, 27] was solved instead of the inviscid equation, in which the rotation of the vortices can be considered. In contrast, (iv) we assume spin angular momentum conservation of fluid particles around their axes; thus, the viscosity term is decomposed into three terms in Eq. (2). The aforementioned (i) corresponds to Figs. 1(a) and (b), (ii) corresponds to (e), (iii) corresponds to (c), and (iv) corresponds to (d). It should be emphasized that this paper is the first to clearly describe that cryogenic liquid helium-4 is a system in which viscous fluid components “remain” in the inviscid fluid at finite temperatures, and that the residual components decrease with decreasing temperature, becoming a “uniform inviscid fluid” at absolute zero.
Let us summarize the physical significance of each parameter in Eqs. (1) and (2) with reference to the literature [7]. $\rho_{n}$ and $\rho_{s}$ are the mass densities of the normal and inviscid fluid components, respectively, which satisfy the relationship $\rho=\rho_{n}+\rho_{s}$, where $\rho$ is the mass density of the total fluid. $D\{\cdot\}/Dt$ denotes the material derivative. $\bm{\mathbf{v}}_{n}$ and $\bm{\mathbf{v}}_{s}$ are the velocities of normal and inviscid fluid components, respectively. $P$, $T$, and $\sigma$ denote the pressure, temperature, and entropy density, respectively. $\bar{\eta}$, $\bar{\xi}$, and $\bar{\eta_{r}}$ indicate the shear, bulk, and rotational viscosities, respectively, where $\bar{\eta_{r}}$ determines the impact of the rotational forces [10]. The rotational force here indicates the rotational force around the axis of each fluid particle. The vector $\bm{\mathbf{\omega}}_{0}$ in the fifth term on the right-hand side of Eq. (2) represents the angular velocity of each fluid particle along its axis. While the vorticity is defined on a continuum, the vector $\bm{\mathbf{\omega}}_{0}$ is a variable of the microscopic constituent particles, which in its original formulation represents molecular particles and later approximates fluid particles. Hence, the vorticity and $\bm{\mathbf{\omega}}_{0}$ are different variables. This is explained further. Originally, Condiff [9] reformulated the Navier–Stokes equations from Cauchy’s equation of motion and the conservation law of angular momentum to include the internal freedom of molecular particles to account for the effect of their spins on the entire polar fluid, resulting in a Navier–Stokes equation with spin-angular momentum conservation. Because molecules are polarized in polar fluids, the influence of the rotational motion of individual molecules on the bulk fluid needs to be considered. A schematic of the derivation of this equation is presented in the Appendix Fig. 9. It is worth noting that Condiff divided the total angular momentum ($\bm{\mathbf{M}}$) per unit mass of the fluid into the angular momentum of the bulk fluid ($\bm{\mathbf{r}}\times\bm{\mathbf{u}}$) and the contribution due to the internal degrees of freedom of the molecules, that is, the rotational angular momentum ($\bm{\mathbf{I}}$) (see Fig. 9(D)). The vectors $\bm{\mathbf{r}}$ and $\bm{\mathbf{u}}$ represent the coordinates and velocity of the local fluid fragment, respectively. $\bm{\mathbf{I}}$ can be further decomposed into $\bm{\mathbf{\it I}}\cdot\bm{\mathbf{\omega}}_{0}$, where $\bm{\mathbf{\it I}}$ is a tensor field, which is a scalar multiple of the unit dyadic if we postulate a uniform and isotropic rotational field (and thus a rigid-body rotation) for each constituent particle. In this context, $\bm{\mathbf{\omega}}_{0}$ denotes the angular velocity of a molecule’s rotation around its axis. Thus, $\bm{\mathbf{\omega}}_{0}$ is in principle a quantity defined for a molecule. However, $\bm{\mathbf{\omega}}_{0}$ is often redefined for each fluid particle comprising a collection of molecules, assuming that each fluid particle is a representative point among the subordinate molecules based on a coarse-grained approximation. Therefore, the term “fluid particle” here is closer to the definition in the context of molecular fluid dynamics, and thus, may be not the same as a fluid particle as defined in classical fluid mechanics. When considering mesoscale flow problems, such as molecular fluids, fluid particles are often interpreted in both the classical and microscopic senses; we still refer to them as classical to compare the quantum mechanical description of the ordinary two-fluid model in this paper.
In the Lagrangian picture, the difference between the vorticity ×$\nabla\times\bm{\mathbf{u}}$ owing to the rotational motion of the bulk fluid and the vorticity owing to the rotation of the molecules can be assumed to cause strain. Specifically, we assume that the deviation of the vorticity from the rigid-body rotation yields strain because the molecules rotate rigidly. When the vorticity ×$\nabla\times\bm{\mathbf{u}}$ at a point is a rigid-body rotation, the local rotation of the fluid is synchronized with the spin of the molecule; thus, ×$\nabla\times\bm{\mathbf{u}}-2\bm{\mathbf{\omega}}_{0}=0$. In this case, no viscous stress owing to local rotation occurs. Otherwise, the difference in ×$\nabla\times\bm{\mathbf{u}}-2\bm{\mathbf{\omega}}_{0}$ induces strain in the fluid, making it that deviate from rigid-body rotation, and contributes to the asymmetric part of the stress tensor (see Figs. 9(F) and (I) in the Appendix). In the third through fifth terms on the right-hand side of Eq. (2), the viscosity term is divided into three contributions: shear, volume, and rotational viscosities. Thus, the spin-angular momentum-conserving Navier–Stokes equations can be rephrased as Navier–Stokes equations with reformulated viscosity terms. This model can easily connect microscopic and fluid mechanics in a continuum mechanical regime. To date, it has been applied to mesoscale flows in which the rotation of molecules or fluid particles dominates, such as in polar fluids and suspension flows. Previously, Müller adopted SPH to discretize the spin-angular momentum-conserving Navier–Stokes equations to simulate two immiscible flow [10]; their model was further applied to red blood cell flow [28].
Mathematically, SPH is a finite particle approximation of a continuum field that can decompose the quantities defined in the field into neighboring particle contributions. The intrinsic part of SPH is a representation of a physical quantity in an integral form using the Dirac delta function $\delta$ and its approximation by the finite distribution function $W$, called the smoothed kernel function. We then discretize the approximation form based on the concept of the sum approximation for numerical simulations. This paper does not focus on the derivation of the SPH models because they have been discussed in our previous studies [6, 7, 8]; however, the basic concepts of SPH and its discretized models used in this study are illustrated in Fig. 10 in the Appendix. In Refs. [6, 7, 8], the viscosity term of the two-fluid model for cryogenic liquid helium-4 was reformulated to conserve the spin-angular momentum. In other words, the viscosity term was reformulated so that the two-fluid model can be applied not only to shear viscosity-dominated problems such as countercurrent flow, but also to rotation-dominated problems. We then discretized the equations using the SPH. Consequently, we observed the formation of multiple spinning vortices that occurred independently of the overall fluid motion in our SPH simulations of the rotating 2D liquid helium-4 [6]. After incorporating the vortex dynamics model into our model, we found that multiple spinning vortices formed a lattice with a regular arrangement [7, 8]. The number of vortices generated was determined based on the strength of the rotational angular velocity. For appropriate values, the number of vortices agreed well with the theoretical solution of the number of vortices calculated using Feynman’s law [7]. Compared with our previous studies, this study is the first to address a 3D problem. The system does not assume gravity in any direction, and considering numerical stability, the validation in the 2D cases is effective for the 3D cases as well. As we focus on the disk-shaped problem and consider only the horizontal rotation around the vertical axis (z) of the disk, the developed rotational viscosity term for the two-dimensional model can be used directly. However, because of the topological differences in vortices between the 2D and 3D cases, it needs to be studied whether the previous method, which is effective in solving the 2D rotating problem, can numerically reproduce the vortices in 3D liquid helium-4; large-scale simulations for 3D cases will be reported and discussed later.
@@
Let us explain the physical significance of each parameter in Eqs. (3) and (4) with reference to Refs. [6, 7]. In Eq. (3), $M$ represents the mass of a helium-4 atom, $N$ indicates the total number of helium-4 atoms in the system, $\pi$ represents the ratio of the circumference of a circle to its diameter, and $c$ is the speed of sound. $\hbar$ indicates the reduced Planck’s constant and $k_{B}$ represents the Boltzmann constant. The constant values $\mu$, $p_{0}$, and $\Delta$ are given by ${\rm 1.72\times 10^{-24}~{}g}$, ${\rm 2.1\times 10^{-19}~{}gcms^{-1}}$, and ${\rm 8.9~{}K}$, respectively  [31, 32]. Eq. (3) is always true within the scope of this study because it is valid if $T\ll{\rm 93~{}K}$ [31]. In the elementary excitation model, the first term in the parentheses on the right side of Eq. (3) represents the “photon” contribution, and the second term in the same parentheses represents the “roton” contribution; after finding the total pressure from the summation of the photon and roton contributions, the entropy density in Eq. (3) is obtained by the thermodynamic relation, such that the partial derivative of pressure with respect to temperature [6]. Equation (3) represents the relationship between entropy density $\sigma$ and temperature $T$ in a quantum “ideal gas.” To obtain the value of liquid helium-4, we decompose $\sigma$ as $\sigma=C_{e}\sigma_{0}$, where $C_{e}$ indicates the volume ratio of liquid to gaseous helium-4 under an atmospheric pressure of $\rm 1.01325~{}bar$, which is $1.428\times 10^{-3}$ [33]. Under constant temperature, we calculated $\sigma_{0}$ using Eq. (3) and maintained a constant value for entropy conservation. Meanwhile, in Eq. (4), $\rho_{0}$ and $p_{0}$ represent the fluid density and pressure in the initial state, respectively. $\alpha_{P}$ is the specific heat ratio that affects system incompressibility. $\beta_{P}$ denotes the reference pressure, which typically corresponds to $p_{0}$. Equations (1)–(4) represent the governing equations of the system.
2.2 The potential and immediate challenges of our two-fluid model
Notably, for the dynamics of cryogenic liquid helium-4, both the derivation of the governing equations from the mechanical interpretation in Figs. 1(a)–(e) and quantum mechanics resulted in similar mathematical forms. The first model assumes a two-phase flow model in which the two components are separated and the fluid particles have a volume exclusion effect on each other. Particle correction has been introduced as an exception. Conversely, the second or ordinary two-fluid model assumes that both components coexist in space and that neither type of particle exerts volume exclusion effects on each other, and introduces mutual friction as an exception. In summary, the governing equation can be identical regardless of whether the system comprises separate or coexisting (overlapping) particles. However, separate and overlapping constituent particles lead to different physical meanings, and the former and latter models provide the classical and quantum-mechanical descriptions of the same problem. In particular, the physical meaning of “particles” is different; the former indicates classical fluid particles, and the latter represents quantum minute particles. This inspired us to reconceptualize the system in terms of the degree of volume exclusion between the two components, hereafter, referred to as the mutual volume exclusion (MVE) effect. Here, the MVE effect can be rephrased as the effect of classical fluid forces, such as surface forces, body forces, or interface tensions. This eliminates the need to distinguish the physical meanings of the particles between the two models, considering that the MVE is the primary factor causing the particles to separate or overlap.
Ultimately, the two models can be unified into one model in terms of the MVE effect as a variable. This model has the following advantages. New insights into the connection between classical and quantum mechanics can be obtained by using the MVE effect as a unified metric of the system. This may also provide a new perspective on the relationship between the classical two-phase flow and the ordinary two-fluid model of unified hydrodynamics. In engineering and industry, it may become possible to capture classical and quantum fluid phenomena with only one model by adjusting the degree of MVE of the constituent particles depending on the location and condition. Practically, there is no need to switch computational codes depending on the conditions, and a single computational code can be used to analyze the entire large-scale system.
In such a unified model, the transition from classical to quantum particles can, in principle, be automatically performed using high-resolution computational calculations. This is because as the computational resolution increases, the size of the fluid particles in the classical domain gradually decreases, and the excluded volume effect of individual fluid particles also decreases. In other words, a low computational resolution in a local region implies that the size of the analytic particles at that location is large, and a large analytic particle size implies that the excluded volume effect of the individual particles is large, which is suitable for describing the location where the classical fluid is dominant. Conversely, a high computational resolution in a local region implies that the size of the analytical particles at that location is small, which implies that the excluded volume effect of individual particles is small, which is appropriate for a local region where quantum hydrodynamics dominate. In summary, in the unified model, the convergence of the computational resolution is expected to correspond to the convergence of the physical phenomena. In this case, the introduction of multiresolution particles makes it possible to reproduce the behavior of the fluid in the entire system in a single model.
To this end, the immediate challenge of our approach is to provide evidence that increasing the resolution brings the simulation results closer to the phenomena observed in the quantum-mechanical regime, even when using our classical mechanical two-fluid model with spin-angular momentum conservation.
@@ 2.3 Multi-GPU computing for large-scale particle simulations
Figure 2: Snapshots of (a) the particle distribution, with inviscid and viscid fluid particles colored in red and blue, respectively, and (b) the domain decomposition in the case where approximately 19.6 million fluid particles were divided into 32 subdomains, each assigned to a single GPU.
It is imperative to develop dynamic load balancing among the GPUs to exploit the high computational efficiency of multiple GPUs. Various load balancing techniques have been proposed in related fields, including the slice-grid method, which moves the domain boundary to a spatially partitioned region; graph theory-based partitioning (such as k-means), as represented by ParMETIS [56, 57] and Zoltan [58, 59]; and the hierarchically structured tree method, which uses a recursive structure of cells to partition a domain, as exemplified by kd-tree decomposition [60], orthogonal recursive bisection (ORB) [61, 62, 63], and domain decomposition using a hierarchically structured grid with space-filling curves [64]. Recently, we conducted a study on optimal dynamic domain decomposition for particle simulations of short-range interactions compatible with low byte/flops supercomputers, such as GPU supercomputers. Subsequently, we developed a multi-GPU computing framework that enables large-scale particle simulations using SPH. The SPH computational framework [52] developed by the author was used in this study. Based on the nature of the target problem, the 2D slice grid method was used as the domain decomposition method. Figure 2 illustrates a snapshot of the domain decomposition in the case where the 19.6 million fluid particles were divided into 32 subdomains, each of which was assigned to a single GPU to speed up the simulations.
@@
Figure 3: Geometrical setup of a horizontally rotating 3D liquid helium-4: (a) top view, (b) cross section at y = 0, and (c) bird’s eye view.
@@ 3 Large-scale simulations of horizontally rotating 3D liquid helium-4
Let us set the outer diameter of a cylindrical container in the horizontal direction to 0.2 cm with a thickness of 0.04 cm in the vertical direction (z-direction). The resolutions of $(N_{x},N_{y},N_{z})$ are given by (500, 500, 100), where $N_{a}$ $(a=x,y,z)$ indicates the number of particles in the direction in a rectangular domain that covers the cylindrical container. We set the rotational angular velocity around the cylinder axis (z axis) to 5 rad $\cdot s^{-1}$ counterclockwise. Similar to the two-dimensional cases in our previous work, we set the temperature T to 1.6 K and determined the values of $(\rho_{s}/\rho,\rho_{n}/\rho,\kappa)$ with those at T = 1.6 K listed in Ref. [69]. In addition, $\eta_{n}$ was obtained from the relationship of 0.566 $\kappa\rho_{n}$ with reference to [69]. The shear viscosity $\bar{\eta}$ and rotational viscosity $\bar{\eta_{r}}$ have the value of $\eta_{n}$. We set the parameter $C_{\omega}$, which determines the magnitude of the angular velocity $\bm{\mathbf{\omega}}_{0}$ and thereby determines the resulting number of vortices, to be 0.016 to maintain a similar degree of magnitude of the rotational forces as those in 2D cases reported in our previous studies. Normal fluid particles were placed along the inner circumference of the container as walls, and no-slip conditions were imposed on the wall particles as boundary conditions. Conversely, we set slip conditions at the top and bottom of the vessel. We applied a well-established wall and dummy particle method [70, 71]; we update the density and pressure of the inner wall particles similarly to fluid particles, and retain the initial pressure and density values of the outer wall particles. Figure 3 illustrates the geometrical setup of the target system. In the initial state, superfluid or normal fluid particles were randomly generated in proportion to the density ratio and distributed in the vessel.
The computational conditions in this study were adopted from the parameter values determined in our previous study [6]. In addition to satisfying the Courant-Friedrichs-Lewy (CFL) condition, several benchmark problems were solved to validate the stability and numerical accuracy of the calculations. Specifically, we simulated the propagation of the wave equation using the proposed SPH method and compared its numerical solution with the numerical and theoretical solutions obtained using the difference method. In addition, we compared the simulation results with the exact solution for the case in which the Reynolds number $R$ of the normal flow component was 100, and for the case in which $R$ was 1000 in 2D driven flow problems. We obtained very good agreement between the simulation results and the exact solutions obtained by the high-order FDM [72]. Furthermore, assuming an ideal case in which the two components behave as perfect classical fluids, we solved the Rayleigh-Taylor instability problem under conditions where the density difference $(\rho_{s}/\rho_{n})$ was five times, which is the same as that for the problem covered in this study, and confirmed that the characteristic mushroom structure could be reproduced. These results were confirmed for 2D problems. Although we targeted a 3D problem, we did not need to impose gravity on the system. In addition, the target problem was quasi-2D rather than symmetric 3D, and dynamic changes in the vertical direction were not dominant. Accordingly, we used the values used in previous simulations. Note that we set $\delta t$ to be sufficiently small to satisfy the CFL condition for all resolution cases. We set $\delta t$ to $5.0\times 10^{-6}$ s and computed 180000 time steps, which is 0.9 s in physical time. In the simulation, under the aforementioned conditions, the total number of normal and inviscid fluid particles, including the wall particles, was 19636400. We boosted our simulation using 32 GPUs installed on the Wisteria/BDEC-01 supercomputer at The University of Tokyo, JAPAN. We applied a two-dimensional slice-grid technique to the dynamic domain decomposition in the horizontal direction and decomposed the computational domain into 32 subdomains (four in the horizontal direction and eight in the depth direction), and a single GPU was assigned to each subdomain. Consequently, approximately 5.5 days were required to compute all 180000 time steps of the simulation. For comparison, we performed simulations with a lower resolution using $(N_{x},N_{y},N_{z})$ = (250, 250, 50) to investigate the dependence of the observed phenomena on the computational resolution.